OpenStudy (annie96):

What is the slope of a line parallel to the line whose equation is 2x-3y=6? 1/5 2/3 4/3 3/3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Recall the slope-intercept form: $$\Huge y=mx+b$$

OpenStudy (annie96):

4/3

OpenStudy (skullpatrol):

$$\Huge 2x-3y=6$$

OpenStudy (skullpatrol):

$$\Huge-3y=-2x+6$$

OpenStudy (skullpatrol):

$$\Huge y=\frac{-1}{3}(-2x+6)$$

OpenStudy (annie96):

So the answer will be 2/3.

OpenStudy (skullpatrol):

You are correct :)

OpenStudy (annie96):

thank you so much.

OpenStudy (skullpatrol):

Thanks for trying to learn.